OpenStudy (anonymous):

Compare. Write =, less than, or greater than. 1. 1 over 7 compared to 1 over 8 2. 3 over 4 compared to 11 over 16

OpenStudy (anonymous):

1/7 > 1/8 3/4 > 11/16

OpenStudy (anonymous):

There are lots of different ways of solving a problem like this, but I'll show you one that works well, using #2. Before we start, we don't know whether 3/4 is less than or greater than 11/16, so it doesn't seem like we can write an inequality. But, we can start by writing an inequality and just not deciding whether the sign is < or >: \[\frac{3}{4}~ ? ~\frac{11}{16}\] Then we can cross multiply, just like we would with a standard inequality: \[3 \times 16 ~?~ 4 \times 11\] \[48 ~?~ 44\] And now, it's obvious that ? should be <. So, that's the answer: 3/4 is less than 11/16.

OpenStudy (texaschic101):

or you could turn them into decimals....or you can find the common denominator and work it that way. There are a few ways, you just need to pick the best way that works for you.